# Microphone Array Geometry Independent Multi-Talker Distant ASR: NTT System for the DASR Task of the CHiME-8 Challenge

### v1

@@ 3.3 Multi-channel integration with DOVER-Lap
Various approaches have been proposed to integrate the results from multiple microphones. One possible approach is the early fusion approach, which leverages fine spatial information, like spatial feature clustering [48] or incorporating multi-channel features to the input of EEND[34], both of which could offer valuable insights for speaker discrimination. However, such approaches may struggle with tracking speakers who move frequently and could be sensitive to the layout of the microphone array. An alternative is the late fusion approach, which combines diarization results from each microphone independently. While the late fusion does not directly access the fine spatial details, its independent handling of each channel may offer greater robustness to speaker movement and variations in an array configuration.
In our proposed system, we adopt the latter approach, performing diarization independently on each channel using EEND-VC. Specifically, we merge the diarization results from all available microphones with DOVER-Lap [33]. For DOVER-Lap, we employ the Hungarian algorithm to find the optimal speaker permutation between the diarization results across different microphones and integrate them using a voting scheme. This voting scheme can help mitigate the impact of outlier results, such as those from microphones far from the sources or affected by recording failures. The effectiveness of DOVER-Lap-based microphone integration will be demonstrated in an ablation study in A.1.3
@@ 3.4 TS-VAD refinement (DIA2/DIA3)
After obtaining the initial diarization results using the EEND-VC module, we refine these results with target speaker voice activity detection (TS-VAD). Specifically, we adopt the sequence-to-sequence TS-VAD (Seq2Seq-TS-VAD) [27], applying it separately to each observed microphone signal to obtain microphone-dependent global speaker activity, as illustrated in Fig. 5.
@@
For the encoder, we used Conformers consisting of a convolutional NN (CNN)-based downsampler, followed by six Conformer blocks to transform the audio into frame-wise features. For the decoder, we employ a six-layer Transformer-based structure equipped with a memory-aware multi-speaker embedding (MA-MSE) module [27]. Session-level i-vectors are employed as session-wise speaker embeddings, obtained by averaging the segment-level i-vectors belonging to the same speaker. It then combines these i-vectors with local speaker embeddings derived from the input mixture and the local segmentation information, which is implicitly performed within the speaker-wise decoder block in Fig 4. These combined embeddings are used to condition a Conformer-based TS-VAD module. Our system has almost the same model configuration as that used in the top system of CHiME-7 [17] but with a stronger initial diarization provided by EEND-VC. We confirm the importance of the initialization with EEND-VC in A.1.1.
We obtain the diarization results from the speaker activity posteriors using the following post-processing steps. First, the global speaker activities are stitched together and threshold using $\theta_{\textrm{ts-vad}}$. The boundaries of each utterance are then padded by a specified offset, $\tau_{\mathrm{offset}}$, and short pauses shorter than $\tau_{\mathrm{merge}}$ are filled with ones. Figure 6 illustrates how these three parameters work. These processes prevent the occurrence of spurious utterances while avoiding the omission of utterance boundaries and pauses, enabling better segmentation results for downstream ASR modules. These steps are performed on the activities obtained for each microphone. These activities are subsequently integrated into microphone-independent global speaker activities using DOVER-Lap, as in the pre-diarization step described in Section 3.3. Note that boundary padding and short pause filling are only applied for ASR inputs, and the boundaries obtained by thresholding alone are used for GSS.
@@
Figure 6: Postprocessing to obtain utterance boundaries from speaker activity. The blue plot represents the original activity from a speaker, the green plot shows the result after thresholding using $\theta_{\mathrm{ts-vad}}$, the purple plot represents the result after padding by an offset, $\tau_{\mathrm{offset}}$, and finally, the red plot illustrates the activity obtained after filling pauses of duration $d$ shorter than $\tau_{\mathrm{merge}}$
@@ 3.5.1 Model configuration
EEND: We utilized the pre-trained WavLM-large model as a feature extractor, following the approach used in [49]. Outputs from all WavLM Transformer layers were averaged using learnable weights, resulting in 1024-dimensional input speech features. These features were then fed into the EEND module, following the same configurations as in our previous CHiME-7 submission [19]. The EEND module consisted of six stacked Transformer encoder blocks with eight 256-dimensional attention heads, followed by a linear layer that projected the encoder’s output into four streams, representing the frame-by-frame speaker activity probabilities for the four local speakers (i.e., $S_{\mathrm{local}}=4$). We applied thresholds of $\theta_{\mathrm{emb}}=0.5$ and $\theta_{\mathrm{eend}}=0.5$ to the outputs of the EEND modules to extract ECAPA-TDNN embeddings, as described in Eq. (10), and to obtain binarized speaker activity during the post-processing step, respectively.
Vector clustering: We used ECAPA-TDNN [41] for speaker embedding extraction, applying it to 30-second chunks ($T_{\mathrm{c}}=30$) that were enhanced using a chunk-wise SE module. We employed SpeechBrain’s ECAPA-TDNN [50], pre-trained on over 2000 hours of audio from more than 7000 speakers in VoxCeleb1 and VoxCeleb2 [51]. Subsequently, constrained spectral clustering was performed on these speaker embeddings, with the speaker count estimated using the speaker counting module.
@@
TS-VAD: The i-vector extractor was trained on VoxCeleb1 and VoxCeleb2 using the Kaldi recipe from egs/voxceleb/v1444https://github.com/kaldi-asr/kaldi/tree/master/egs/voxceleb/v1. We computed 40-dimensional Mel-frequency cepstral coefficients (MFCCs) calculated with a 25 ms window size and a 10 ms shift, followed by mean normalization using three-second sliding windows as acoustic features for i-vector extraction. Silence regions and utterances shorter than 100 ms were removed based on local speaker activity estimated by a pre-diarization module. Subsequently, i-vectors were extracted for each speaker’s utterance and averaged across the session to generate session-wise i-vectors for each speaker. The implementation of Seq2Seq-TS-VAD relied on the publicly available implementation of NSD-MS2S [27, 52]. The input to NSD-MS2S consisted of 40-dimensional filterbank features computed using a 25 ms window size and a 10 ms shift, followed by mean normalization with three-second sliding windows. We used the default parameters implemented in NSD-MS2S code [^5] https://github.com/liyunlongaaa/NSD-MS2S, except that we used only a single deep interactive module block.
Post-processing: To allow faster experimental turnaround, we tuned most parameters of the diarization module to achieve the best diarization error rate (DER) on the development set. However, DER is not well correlated with recognition performance. Therefore, after the initial tuning, we performed an extensive grid search on the optimal post-processing parameters after TS-VAD for ASR, including the activity threshold, boundary offset, and short-pause filling durations. As a result of the tuning, we obtained $\theta_{\textrm{ts-vad}}=0.30$ , $\tau_{\mathrm{offset}}=0.0$ sec, and $\tau_{\mathrm{merge}}=1.5$ sec as the best parameters. The effect of this tuning is reported in A.1.4.
@@ 3.5.3 Training procedure
TS-VAD: For both the initial training and fine-tuning, we optimized all NSD-MS2S parameters to minimize the binary cross entropy between the model output and the ground truth labels. We pretrained the model for six epochs on DIA2 and three epochs on DIA3, respectively. We used Adam optimizer with a learning rate of $10^{-3}$ and a batch size of 256. For fine-tuning, we again updated all NSD-MS2S parameters using the CHiME-8 training set for six epochs with a learning rate of $10^{-5}$ and a batch size of 1024. Following the procedure used in [27], we averaged the model parameters obtained over six epochs in the fine-tuning step.
@@ 4 Speech enhancement
This section describes how we implement our SE system described in Eq. (4). Figure 7 shows the processing flow of our SE system, which follows the same structure as the official SE system [14] implemented in the CHiME-8 DASR NeMo Baseline [3, 46]. These SE systems rely heavily on Guided Source Separation (GSS) [14] to robustly separate the target speaker with the help of speaker activity provided by the diarization system. They also feature microphone subset selection and beamformer reference microphone selection to effectively handle distributed microphones and maximize their potential. We briefly review the official SE system in Section 4.1 and then describe our two key modifications to it in Sections 4.2 and 4.3.
Figure 7: Processing flow of our SE system, which follows the same structure as the CHiME-8 DASR NeMo baseline system. While the baseline system uses R1-MWF for the MIMO beamforming module, our SE system uses SP-MWF (Section 4.3). In addition, our microphone subset selection module differs from that of the baseline system (Section 4.2).
Figure 8: Processing flow of the SE system used for speaker diarization. Unlike the SE system, signals from all microphones are utilized for WPE and beamforming. Additionally, GSS is applied based on local activity estimated on each of the $M^{\prime}$ channels by the EEND module.
4.1 The CHiME-8 DASR NeMo baseline
@@
In a distributed microphone scenario, some microphones may have a significantly lower signal-to-noise ratio (SNR) than others, and such microphones may not help or may even negatively affect SE. To ignore such uninformative microphones, the microphone subset selection module in Figure 7 selects the top $K$ microphones among all $M$ microphones that are considered to be the most relevant to the subsequent SE processing. Selecting fewer microphones also helps reducing the computational load of the SE system. To select the top $K$ microphones, the baseline system borrows a microphone ranking method from [59], which is based on an audio feature called envelop variance (EV). See Section 4.2 for more details. Note that the microphone subset selection is carried out independently for each utterance segment $\mathbf{b}_{s,j}=[t^{\text{start}}_{s,j},t^{\text{end}}_{s,j}]$.
@@
After selecting the top $K$ microphones for segment $\mathbf{b}_{s,j}=[t^{\text{start}}_{s,j},t^{\text{end}}_{s,j}]$, the baseline system expands the segment to $\mathbf{b}_{s,j}^{\Delta}=[t^{\text{start}}_{s,j}-\Delta,t^{\text{end}}_{s,j}+\Delta]$, where $\Delta\geq 0$ is said to be a context length and set to $15$ seconds in the default setting. Expanding the segment is primarily to enable the GSS module to estimate the time-frequency (TF) masks accurately. After selecting $K$ microphones, we cut out the $K$-channel signal with expanded segment $\mathbf{b}_{s,j}^{\Delta}$, transform the cut signal to the short-term Fourier transform (STFT) domain, and apply dereverberation to it. For dereverberation, the baseline system adopts a multiple-input multiple-output (MIMO) implementation [28] of weighted prediction error (WPE) [29], which has served as an essential SE component in the past CHiME Challenge series.
@@
The dereverberated $K$-channel signal with expanded segment $\mathbf{b}_{s,j}^{\Delta}$ is passed to the GSS module to estimate a single channel time-frequency (TF) mask for separating target speaker $s$ on segment $\mathbf{b}_{s,j}^{\Delta}$. GSS [14] is a clustering-based TF mask estimation method, building on a blind method called the complex angular central Gaussian mixture model (cACGMM) [15]. GSS extends cACGMM to be able to exploit the speaker activity of all sources, i.e., $\hat{\mathbf{B}}$ in Eq. (4), which is provided by the diarization system. Guided by the speaker activity, GSS can estimate the TF masks for all sources more accurately than cACGMM and further associate each TF mask with the corresponding source, which is crucial for separating a target speaker $s$. For more information on GSS, see [14].
@@
After GSS estimates the TF masks, the beamforming module cuts off the added context in $\mathbf{b}_{s,j}^{\Delta}$ from the $K$-channel dereverberated signal and optimizes an arbitrary MIMO beamformer for separating target speaker $s$ over original target segment $\mathbf{b}_{s,j}$ (not over $\mathbf{b}_{s,j}^{\Delta}$) assuming that the $j$th utterance of target speaker $s$ is correctly included in segment $\mathbf{b}_{s,j}$. The baseline system uses an effective implementation of the minimum variance distortionless response (MVDR) beamformer [30, 60], which is sometimes called Souden MVDR or the rank-1 multichannel Wiener filter (R1-MWF). This beamformer avoids error-prone estimation of the target source steering vector and has been empirically proven effective when used as frontend SE systems for ASR through past CHiME challenges. We describe in Section 4.3 how we obtain the MIMO beamformer using the TF masks.
@@
The MIMO beamformer has $K$ separation filters, or $K$ separated signals, each designed to estimate a target clean signal as observed at each microphone. The reference microphone selection module selects, among $K$ separated signals, the separated signal or the corresponding reference microphone that is considered to be the most effective in some sense. For this purpose, the baseline system is based on the criterion of maximizing the output SNR of the beamforming output signal [61, 62]. See Section 4.3.4 for a detailed description.
@@
Finally, given the single-channel beamforming output signal, the post-filtering module applies the blind analytic normalization (BAN) [63] post-filter to it for adjusting the frequency-wise amplitude of the enhanced signal. After that, using the TF mask for the target speaker provided by the GSS module, the post-filtering module also applies a TF masking to further enhance the signal. We will further discuss this post-filtering module in Section 4.3.5.

### v2

@@ 3.3 Multi-channel integration with DOVER-Lap
Various approaches have been proposed to integrate the results from multiple microphones, including early fusion [34, 48] and late fusion [33, 49]. The early fusion approach leverages fine spatial information, by applying muti-channel speech enhancement before diarization [50, 51], spatial feature clustering [48], or incorporating multi-channel features to the input of EEND [34], all of which could offer valuable insights for speaker discrimination. However, such methods may struggle with tracking speakers who move frequently and might be sensitive to the layout of the microphone array. An alternative is the late fusion approach, which combines diarization results from each microphone independently. We can perform late fusion by using techniques such as posterior combination [49] or DOVER-Lap [33]. While the late fusion does not directly access the fine spatial details, its independent handling of each channel may offer greater robustness to speaker movement and variations in an array configuration.
In the proposed system, motivated by prior studies [33] and our preliminary experiments, we adopt the late fusion approach using DOVER-LAP to merge the diarization results from all available microphones. For DOVER-Lap, we utilize the Hungarian algorithm to find the optimal speaker permutation between the diarization results across different microphones and integrate them using a voting scheme. Specifically, we first aligns clustering labels across channels using the Hungarian algorithm and then aggregates the results using majority voting. This voting scheme can help mitigate the impact of outlier results, such as those from microphones far from the sources or affected by recording failures. The effectiveness of DOVER-Lap-based microphone integration is demonstrated in the ablation study in A.1.3.
@@ 3.4 TS-VAD refinement (DIA2/DIA3)
After obtaining the initial diarization results using the EEND-VC module, we refine these results with target speaker voice activity detection (TS-VAD). Specifically, we adopt the sequence-to-sequence TS-VAD (Seq2Seq-TS-VAD) [27], applying it separately to each observed microphone signal to obtain microphone-dependent global speaker activity, as illustrated in Figure 5.
@@
For the encoder, we utilize Conformers consisting of a convolutional NN (CNN)-based downsampler followed by six Conformer blocks to transform the audio into frame-wise features. For the decoder, we used a six-layer Transformer-based structure equipped with a memory-aware multi-speaker embedding (MA-MSE) module [27]. Session-level i-vectors are utilized as session-wise speaker embeddings, obtained by averaging the segment-level i-vectors belonging to the same speaker. These i-vectors are then combined with local speaker embeddings derived from the input mixture and the local segmentation information, which is implicitly performed within the speaker-wise decoder block in Figure 5. These combined embeddings are then used to condition the Conformer-based TS-VAD module. Our system has almost the same model configuration as that in the top system of CHiME-7 [17] but with a stronger initial diarization provided by EEND-VC. We confirm the importance of the initialization with EEND-VC in A.1.1.
We obtain the diarization results from the speaker activity posteriors using the following post-processing steps. First, the global speaker activities are stitched together and thresholded using $\theta_{\textrm{ts-vad}}$. The boundaries of each utterance are then padded by a specified offset, $\tau_{\mathrm{offset}}$, and short pauses shorter than $\tau_{\mathrm{merge}}$ are filled with ones. As show in figure 6, these processes prevent the occurrence of spurious utterances while avoiding the omission of utterance boundaries and pauses, enabling better segmentation results for downstream ASR modules. These steps are performed on the activities obtained for each microphone and the activities are subsequently integrated into microphone-independent global speaker activities using DOVER-Lap, as in the pre-diarization step described in Section 3.3. Note that boundary padding and short pause filling are only applied for ASR inputs, and the boundaries obtained by thresholding alone are used for GSS.
@@
Figure 6: Postprocessing to obtain utterance boundaries from speaker activity. The blue plot represents the original activity from a speaker, the green plot shows the result after thresholding using $\theta_{\mathrm{ts-vad}}$, the purple plot represents the result after padding by an offset, $\tau_{\mathrm{offset}}$, and the red plot illustrates the activity obtained after filling pauses of duration $d$ shorter than $\tau_{\mathrm{merge}}$
@@ 3.5.1 Model configuration
EEND: We utilized the pre-trained WavLM-large model as a feature extractor, following the approach used in [52]. Outputs from all WavLM Transformer layers were averaged using learnable weights, resulting in 1024-dimensional input speech features. These features were then fed into the EEND module, following the same configurations as in our previous CHiME-7 submission [19]. The EEND module consisted of six stacked Transformer encoder blocks with eight 256-dimensional attention heads, followed by a linear layer that projected the encoder’s output into four streams representing the frame-by-frame speaker activity probabilities for the four local speakers (i.e., $S_{\mathrm{local}}=4$). We applied thresholds of $\theta_{\mathrm{emb}}=0.5$ and $\theta_{\mathrm{eend}}=0.5$ to the outputs of the EEND modules to extract ECAPA-TDNN embeddings, as described in Eq. (10), and to obtain binarized speaker activity during the post-processing step, respectively.
Vector clustering: We used ECAPA-TDNN [41] for speaker embedding extraction, applying it to 30-second chunks ($T_{\mathrm{c}}=30$) that were enhanced using a chunk-wise SE module. We utilized SpeechBrain’s ECAPA-TDNN [53], pre-trained on over 2000 hours of audio from more than 7000 speakers in VoxCeleb1 and VoxCeleb2 [54]. Subsequently, constrained spectral clustering was performed on these speaker embeddings, with the speaker count estimated using the speaker counting module.
@@
TS-VAD: The i-vector extractor was trained on VoxCeleb1 and VoxCeleb2 using the Kaldi recipe from egs/voxceleb/v1444https://github.com/kaldi-asr/kaldi/tree/master/egs/voxceleb/v1. We computed 40-dimensional Mel-frequency cepstral coefficients (MFCCs) calculated with a 25 ms window size and a 10 ms shift, followed by mean normalization using three-second sliding windows as acoustic features for i-vector extraction. Silence regions and utterances shorter than 100 ms were removed based on local speaker activity estimated by the pre-diarization module. Subsequently, i-vectors were extracted for each speaker’s utterance and averaged across the session to generate session-wise i-vectors for each speaker. The implementation of Seq2Seq-TS-VAD relied on the publicly available implementation of NSD-MS2S [27, 55]. The input to NSD-MS2S consisted of 40-dimensional filterbank features computed using a 25 ms window size and a 10 ms shift, followed by mean normalization with three-second sliding windows. We used the default parameters implemented in NSD-MS2S code [^5] https://github.com/liyunlongaaa/NSD-MS2S, except that we used only a single deep interactive module block.
Post-processing: To enable faster experimental turnaround, we tuned most parameters of the diarization module to achieve the best diarization error rate (DER) on the development set. However, DER is not well correlated with recognition performance. Therefore, after the initial tuning, we performed an extensive grid search on the optimal post-processing parameters after TS-VAD for ASR, including the activity threshold, boundary offset, and short-pause filling durations. After the tuning, we obtained $\theta_{\textrm{ts-vad}}=0.30$ , $\tau_{\mathrm{offset}}=0.0$ sec, and $\tau_{\mathrm{merge}}=1.5$ sec as the best parameters. The effect of this tuning is reported in A.1.4.
@@ 3.5.3 Training procedure
TS-VAD: For both the initial training and fine-tuning, we optimized all NSD-MS2S parameters to minimize the binary cross entropy between the model output and the ground truth labels. We pretrained the model for six epochs on DIA2 and three epochs on DIA3. We used Adam optimizer with a learning rate of $10^{-3}$ and a batch size of 256. For fine-tuning, we again updated all NSD-MS2S parameters using the CHiME-8 training set for six epochs with a learning rate of $10^{-5}$ and a batch size of 1024. Following the procedure used in [27], we averaged the model parameters obtained over six epochs in the fine-tuning step.
@@ 4 Speech enhancement
In this section, we explain how we implement our SE system described in Eq. (4). Figure 7 shows the processing flow of our SE system, which follows the same structure as the official SE system [14] implemented in the CHiME-8 DASR NeMo Baseline [3, 46]. These SE systems rely heavily on Guided Source Separation (GSS) [14] to robustly separate the target speaker with the help of speaker activity provided by the diarization system. They also feature microphone subset selection and beamformer reference microphone selection to effectively handle distributed microphones and maximize their potential. We briefly review the official SE system in Section 4.1 and then describe our two key modifications to it in Sections 4.2 and 4.3.
Figure 7: Processing flow of proposed SE system, which follows the same structure as the CHiME-8 DASR NeMo baseline system. While the baseline system uses R1-MWF for the MIMO beamforming module, our SE system utilizes SP-MWF (Section 4.3). In addition, our microphone subset selection module differs from that of the baseline system (Section 4.2).
Figure 8: Processing flow of SE system used for speaker diarization. Unlike the SE system, signals from all microphones are utilized for WPE and beamforming. Additionally, GSS is applied based on local activity estimated on each of the $M^{\prime}$ channels by the EEND module.
4.1 CHiME-8 DASR NeMo baseline
@@
In a distributed microphone scenario, some microphones may have a significantly lower signal-to-noise ratio (SNR) than others, and such microphones may not help or may even negatively affect the SE. To ignore such uninformative microphones, the microphone subset selection module in Figure 7 selects the top $K$ microphones among all $M$ microphones that are considered to be the most relevant to the subsequent SE processing. Selecting fewer microphones also helps to reduce the computational load of the SE system. To select the top $K$ microphones, the baseline system borrows a microphone ranking method from [62], which is based on an audio feature called envelop variance (EV); see Section 4.2 for more details. Note that the microphone subset selection is carried out independently for each utterance segment $\mathbf{b}_{s,j}=[t^{\text{start}}_{s,j},t^{\text{end}}_{s,j}]$.
@@
After selecting the top $K$ microphones for segment $\mathbf{b}_{s,j}=[t^{\text{start}}_{s,j},t^{\text{end}}_{s,j}]$, the baseline system expands the segment to $\mathbf{b}_{s,j}^{\Delta}=[t^{\text{start}}_{s,j}-\Delta,t^{\text{end}}_{s,j}+\Delta]$, where $\Delta\geq 0$ is the context length and is set to $15$ seconds as a default. Expanding the segment is primarily to enable the GSS module to estimate the time-frequency (TF) masks accurately. After selecting $K$ microphones, we cut out the $K$-channel signal with expanded segment $\mathbf{b}_{s,j}^{\Delta}$, transform the cut signal to the short-term Fourier transform (STFT) domain, and apply dereverberation to it. For dereverberation, the baseline system adopts a multiple-input multiple-output (MIMO) implementation [28] of weighted prediction error (WPE) [29], which has served as an essential SE component in the past CHiME Challenge series.
@@
The dereverberated $K$-channel signal with expanded segment $\mathbf{b}_{s,j}^{\Delta}$ is passed to the GSS module to estimate a single channel time-frequency (TF) mask for separating target speaker $s$ on segment $\mathbf{b}_{s,j}^{\Delta}$. GSS [14] is a clustering-based TF mask estimation method built on a blind method called the complex angular central Gaussian mixture model (cACGMM) [15]. GSS extends cACGMM to exploit the speaker activity of all sources, i.e., $\hat{\mathbf{B}}$ in Eq. (4), which is provided by the diarization system. Guided by the speaker activity, GSS can estimate the TF masks for all sources more accurately than cACGMM and further associate each TF mask with the corresponding source, which is crucial for separating a target speaker $s$. For more information on GSS, see [14].
@@
After GSS estimates the TF masks, the beamforming module cuts off the added context in $\mathbf{b}_{s,j}^{\Delta}$ from the $K$-channel dereverberated signal and optimizes an arbitrary MIMO beamformer for separating target speaker $s$ over original target segment $\mathbf{b}_{s,j}$ (not over $\mathbf{b}_{s,j}^{\Delta}$) assuming that the $j$th utterance of target speaker $s$ is correctly included in segment $\mathbf{b}_{s,j}$. The baseline system uses an effective implementation of the minimum variance distortionless response (MVDR) beamformer [30, 63], which is sometimes called Souden MVDR or the rank-1 multichannel Wiener filter (R1-MWF). This beamformer avoids error-prone estimation of the target source steering vector and has been empirically proven effective when used as frontend SE systems for ASR through past CHiME challenges. Section 4.3 explains how we obtain the MIMO beamformer using the TF masks.
@@
The MIMO beamformer has $K$ separation filters, or $K$ separated signals, each designed to estimate a target clean signal as observed at each microphone. The reference microphone selection module selects, among $K$ separated signals, the separated signal or the corresponding reference microphone that is considered the most effective in some sense. For this purpose, the baseline system is based on the criterion of maximizing a certain signal quality measure of the beamforming output signal [64, 65]. See Section 4.3.4 for a detailed description.
@@
Finally, given the single-channel beamforming output signal, the post-filtering module applies the blind analytic normalization (BAN) [66] post-filter to it for adjusting the frequency-wise amplitude of the enhanced signal. After that, using the TF mask for the target speaker provided by the GSS module, the post-filtering module also applies a TF masking to further enhance the signal. This post-filtering module is discussed in Section 4.3.5.